# CangLing-KnowFlow: A Unified Knowledge-and-Flow-fused Agent for Comprehensive Remote Sensing Applications

### v1

@@ 3.2 Expert-Informed Procedural Knowledge Base
The PKB is the foundational component for grounding the agent’s planning capabilities in established scientific methodologies. It addresses the problem of planning hallucination by providing the Orchestrator Agent with a structured source of domain-specific procedural knowledge. The PKB is comprised of two core elements: a formalized Tool Schema and a library of hierarchical Workflow Templates, underpinned by a formal mathematical representation. The concrete structure of these elements is illustrated in Figure 3.
@@ 3.2.1 Mathematical Formulation of Workflows
To represent formally the structure and logic of RS analysis procedures, we model each workflow as a Directed Acyclic Graph (DAG). A workflow $W$ is defined as a tuple $W=(V,E)$, where $V$ is a set of vertices representing tool actions and $E$ is a set of directed edges representing dependencies.
Let $\mathcal{T}=\{t_{1},t_{2},\dots,t_{n}\}$ be the set of all available tools. Each vertex $v_{i}\in V$ corresponds to the invocation of a specific tool $t_{i}\in\mathcal{T}$. The properties of each tool $t_{i}$ are defined by its preconditions, $\text{Pre}(t_{i})$, and its effects, $\text{Eff}(t_{i})$. A directed edge $e_{ij}=(v_{i},v_{j})\in E$ exists from vertex $v_{i}$ to $v_{j}$ if and only if an effect of tool $t_{i}$ satisfies a precondition of tool $t_{j}$, i.e., $\text{Eff}(t_{i})\cap\text{Pre}(t_{j})\neq\emptyset$. This edge signifies a direct dependency, where the output of tool $t_{i}$ is a necessary input for tool $t_{j}$.
The acyclic nature of the graph ensures that a workflow is logically sound and executable. This formal DAG representation is central to the CangLing-KnowFlow framework, as it enables the agent to check the logical consistency of the workflow, identify failure points, and systematically perform dynamic adjustments by manipulating the graph structure. In constrast with existing LLM-based agents, the proposed checking manner follows certain practical rules, such as topological sorting, cycle detection, and computing logics, rather than relying purely on probabilities.
@@ 3.2.2 PDDL Tool Schema
To enable the mathematical formulation, every tool available to the agent is described formally using a schema inspired by the Planning Domain Definition Language (PDDL), as shown in Figure 3b. Each tool $t\in\mathcal{T}$ is represented as an action with a clearly defined set of the following three components:
  • 1.
    Parameters: Defines the input arguments for the tool, including data types and constraints.
  • 2.
    Preconditions ($\text{Pre}(t)$): Logical conditions that must be met before tool execution.
  • 3.
    Effects ($\text{Eff}(t)$): State changes that occur upon tool execution.
We formalize the operational space of the agent, denoted as $\mathbf{\Omega}$, as a structured hierarchical taxonomy rather than a simple collection of functions. This design aligns with the intrinsic multi-level nature of RS tasks. Specifically, $\mathbf{\Omega}$ is organized into two categories: (1) Atomic Operators, which handle low-level data manipulation (e.g., geometric correction, projection); and (2) Semantic Operators, which encapsulate complex analysis tasks (e.g., segmentation, inversion) into high-level function calls. This rigorous taxonomy ensures that the derived tool space $\mathcal{T}\subset\mathbf{\Omega}$ provides the necessary granularity for domain-specific benchmarks. By unifying these heterogeneous functions into a machine-readable library, we establish the fundamental building blocks for constructing our workflow DAGs. We will further detail the concrete instantiation of this toolset, including specific GDAL and deep learning algorithms, in Section 4.1.1.
@@ 3.2.3 Workflow Templates
The primary function of this hierarchical structure is to facilitate efficient, top-down planning. By allowing the agent to retrieve and instantiate high-level templates for common tasks, it reduces the planning search space and ensures adherence to expert-validated procedures. This structural guidance constrains the generative process of the LLM, directing it to produce plans that are both syntactically correct and compliant with standard RS protocols.
@@ 3.3 Dynamic Workflow Adjustment
A key innovation of the CangLing-KnowFlow framework is its ability to adjust dynamically the workflow in response to runtime failures, thereby overcoming the brittleness of static plans. This mechanism is the practical realization of the agent’s policy $\mathcal{A}$ from Eq. (1) under conditions of uncertainty. In RS workflows, the source of this uncertainty could be complicated, such as ambiguities in the user’s prompts (e.g., vague spatial definitions), intrinsic data irregularities (e.g., unexpected cloud cover or sensor noise), and unexpected tool behaviors (e.g., non-convergence of neural networks). When a failure occurs for the abovementioned reasons, the agent must compute a new action $a_{t}$ that deviates from the original plan. This is achieved through a structured process composed of three stages: failure detection, a hierarchical repair strategy, and workflow graph manipulations.
@@ 3.3.1 Failure Detection and Diagnosis
The dynamic adjustment process is initiated when the Dynamic Execution Engine detects a failure at a specific vertex $v_{f}\in V$ of the current workflow $W_{current}$. A failure is defined as either an explicit error returned by a tool or an output that fails to meet a predefined quality metric. Upon detection, the engine halts the execution of dependent downstream nodes and reports the failure state, which becomes part of the history $\mathcal{H}_{t-1}$, to the Orchestrator Agent. The Orchestrator then performs a diagnosis by analyzing this history in conjunction with the preconditions of the failed tool $t_{f}$, aiming to identify the root cause of the failure.
@@ 3.3.2 Hierarchical Repair Strategy
Once a failure is diagnosed, the Orchestrator employs a hierarchical strategy to generate a repair plan. This strategy operationalizes the selection of the next action $a_{t}$ by prioritizing experiential knowledge $\mathcal{K}_{evo}$ over generalized reasoning.
@@
    Tier 2: LLM-driven Creative Repair. If no relevant solution is found in $\mathcal{K}_{evo}$, the agent escalates the problem to the LLM. It provides the LLM with the full context ($\mathcal{H}_{t-1}$, $W_{current}$) to generate a novel repair plan, which then defines the action $a_{t}$.
@@ 3.3.3 Workflow Graph Manipulation
The repair plan, which constitutes the chosen action $a_{t}$, is translated into a set of formal operations on the workflow DAG $W_{current}$. These operations, such as node replacement, insertion, or parameter modification, result in an updated workflow $W^{\prime}_{current}$:
  • 1.
    Node Replacement: Replacing the failed vertex $v_{f}$ with a new vertex $v_{r}$ corresponding to an alternative tool $t_{r}$. All incoming and outgoing edges of $v_{f}$ are remapped to $v_{r}$.
  • 2.
    Node Insertion: Inserting a new vertex $v_{new}$ (e.g., for a data preprocessing step) between a predecessor $v_{p}$ and the failed vertex $v_{f}$. This involves removing the edge $(v_{p},v_{f})$ and adding new edges $(v_{p},v_{new})$ and $(v_{new},v_{f})$.
  • 3.
    Parameter Modification: Retaining the vertex $v_{f}$, but altering the parameters associated with its tool $t_{f}$.
After the graph is manipulated, the Orchestrator validates the logical consistency of the new workflow structure before resuming execution from the point of repair. This structured, multi-stage process enables the agent to robustly and efficiently handle a wide range of runtime challenges.
@@ 3.4 Experience-driven Memory and Evolution
The Evolutionary Memory Module is the component that endows the CangLing-KnowFlow agent with the ability to learn and evolve from its operational experience, addressing the critical learning gap in existing agentic workflows. This is achieved through a post-task analysis and knowledge generalization process that operates on a repository of detailed execution traces.
@@ 3.4.1 Structure of the Evolutionary Memory
The memory module serves as the persistent storage for the evolutionary knowledge $\mathcal{K}_{evo}$ introduced in Eq. (1). Structurally, it is organized as a database of execution traces. Each trace, $\mathcal{E}$, is a comprehensive log of a completed task, containing the initial user request, the initial workflow DAG $W_{init}$, the full sequence of actions taken, all intermediate data and tool outputs, any failures encountered, the dynamic adjustments performed, and the final outcome. This rich, structured dataset serves as the raw material for the agent’s learning processes.
@@ 3.4.2 Knowledge Generalization and Refinement
Success Solidification (Positive Feedback): If a task that involved dynamic adjustments is completed successfully, the final, validated workflow $W_{final}$ is considered a candidate for promotion. The agent analyzes the trace to identify the specific sequence of adjustments that led to success. This successful, dynamically generated workflow is then generalized into a new, reusable template. This template, along with metadata describing the context in which it was effective, is added to the PKB. This process continuously enriches the PKB with novel, field-tested procedures.
Failure Attribution (Negative Feedback): In the event of a terminal task failure, the agent performs a failure attribution analysis on the execution trace $\mathcal{E}$. It identifies the specific failure point $v_{f}$, the state of the workflow at that time, and the diagnosed root cause. From this analysis, it synthesizes a new Pattern-Action Rule, formally denoted as $r\in\mathcal{K}_{evo}$. This rule represents a conditional logic that maps a specific failure pattern to a recommended repair action:
if tool(’YOLOvax’) fails on data(type=’SAR’) with error(’low_contrast’),
then action(insert_tool(’Speckle_Suppression’))
These generated rules are stored in a dedicated section of the memory module and are prioritized during the “Knowledge-driven Repair” tier of the dynamic adjustment process. This creates a negative feedback loop, allowing the agent to learn from its mistakes and provide a robust repair strategy when errors occur, thereby significantly reducing the latency and token costs associated with repetitive LLM querying.
Through the dual mechanisms of solidifying successes into templates and distilling failures into repair rules, the CangLing-KnowFlow agent systematically transforms its episodic experiences into durable, procedural knowledge ($\mathcal{K}_{evo}$), enabling it to evolve and increase its accuracy over time.
@@ 4 Database and Benchmark
To ground our agent’s capabilities and rigorously evaluate its performance, we constructed two key assets: (1) an internal PKB to provide the agent with domain-specific knowledge and procedural guidance, and (2) a comprehensive evaluation suite, headlined by our new KnowFlow-Bench, to assess its workflow orchestration capabilities against expert-defined standards.
@@ 4.1 Procedural Knowledge Base Construction
The PKB is the cornerstone of the CangLing-KnowFlow framework, designed to infuse the agent with expert-level procedural knowledge. This section details the practical construction of the PKB, whose formal structure and mathematical underpinnings were defined in Section 3.2. The construction involved a two-stage process of toolset formalization and high-level workflow curation.
@@ 4.1.1 Toolset Formalization
Deep Learning Tools: We designed a suite of high-level deep learning tools, each representing a primary RS analysis task. These tools abstract away the complexity of specific model architectures, focusing instead on the functional transformation from input to output. The suite includes four main categories: semantic_segmentation, object_detection, change_detection, and remote_sensing_inversion. Each tool is defined by a strict JSON schema that specifies its function, parameters, and input/output requirements, ensuring that they can be reliably interpreted and invoked by the agent. For example, the semantic_segmentation tool requires parameters such as image_path, target_classes, and spatial_resolution_m to execute its pixel-wise classification function.
@@ 4.2.1 The KnowFlow-Bench
KnowFlow-Bench is designed to provide a rigorous, end-to-end evaluation of an agent’s ability to orchestrate complex workflows, a capability not adequately covered by existing benchmarks. For example, benchmarks like ThinkGeo [shabbir2025thinkgeo] excel at evaluating step-by-step tool use within a ReAct framework, while others like GeoCode-Bench [hou2025can] focus specifically on geospatial code generation. However, a dedicated benchmark for assessing an agent’s ability to plan, execute, and dynamically adjust a complete, multi-step workflow from a high-level goal has been absent.
To fill the above void, KnowFlow-Bench is built upon the 162 tasks curated for our PKB. For each task, domain experts annotated two distinct, complete, and executable ground-truth workflows, resulting in a benchmark of 324 instances. Unlike the high-level templates in the PKB, each workflow in KnowFlow-Bench is a concrete sequence of function calls with fully specified parameters. These “gold standard” workflows represent the optimal, or a scientifically valid, solution path given the available toolset, serving as the reference against which the agent’s generated workflows are compared for both procedural correctness and final output accuracy.
@@ 4.2.2 Cross-Benchmark Generalization on ThinkGeo
To evaluate the generalization capabilities of our framework, we also tested the CangLing-KnowFlow agent on the ThinkGeo benchmark [shabbir2025thinkgeo]. ThinkGeo is an agentic benchmark comprising 436 tasks grounded in real-world RS imagery, spanning seven major domains such as urban planning and disaster assessment. It provides a suite of 14 predefined tools and evaluates agents based on their ability to follow a ReAct-style reasoning and execution chain. For our evaluation, we adapted our agent to operate within the ThinkGeo environment. This involved extracting the procedural workflow logic from its ReAct-style annotations and constraining our agent to use only the 14 tools provided by their benchmark. This cross-benchmark evaluation serves to validate the robustness and adaptability of our agent’s core reasoning and planning mechanisms when confronted with a novel set of tasks and tools.

### v2

@@ 3.2 Expert-Informed Procedural Knowledge Base
The PKB is the foundational component for grounding the agent’s planning capabilities upon established scientific methodologies. It addresses the problem of planning hallucination by providing the Orchestrator Agent with a structured source of domain-specific procedural knowledge. The PKB is comprised of two core elements: a formalized Tool Schema and a library of hierarchical Workflow Templates, underpinned by a formal mathematical representation. The concrete structure of these elements is illustrated in Figure 3.
@@ 3.2.1 Mathematical Formulation of Workflows
To represent the structure and logic of RS analysis procedures formally, we model each workflow as a Directed Acyclic Graph (DAG). A workflow $W$ is defined as a tuple $W=(V,E)$, where $V$ is a set of vertices representing tool actions and $E$ is a set of directed edges representing dependencies. Let $\mathcal{T}=\{t_{1},t_{2},\dots,t_{n}\}$ be the set of all available tools. Each vertex $v_{i}\in V$ corresponds to the invocation of a specific tool $t_{i}\in\mathcal{T}$. The properties of each tool $t_{i}$ are mainly defined by its two predefined mapping functions, precondition $\text{Pre}(t_{i})$, and its execution effects, $\text{Eff}(t_{i})$. The set of edges is constructed based on semantic compatibility, with a directed edge $e_{ij}=(v_{i},v_{j})\in E$ exists from vertex $v_{i}$ to $v_{j}$ if and only if an effect of tool $t_{i}$ satisfies a precondition of tool $t_{j}$, defined as
$E=\{e_{ij}|\phi_{LLM}(\text{Eff}(t_{i}),\text{Pre}(t_{j}))=1\},$ (2)
where $\phi_{LLM}$ serves as a boolean indicator function implemented by an LLM to determine if $\text{Eff}(t_{i})$ conceptually entails $\text{Pre}(t_{j})$.
The acyclic nature of the graph ensures that a workflow is logically sound and executable. This formal DAG representation plays a central role in the CangLing-KnowFlow framework, as it enables the agent to check the logical consistency of the workflow, identify failure points, and systematically perform dynamic adjustments by manipulating the graph structure. In contrast with existing LLM-based agents, the proposed checking manner follows certain practical rules, such as topological sorting, cycle detection, and computing logics, rather than relying purely on probabilities.
@@ 3.2.2 PDDL Tool Schema
To enable the mathematical formulation, every tool available to the agent is described formally using a schema inspired by the Planning Domain Definition Language (PDDL), as shown in Figure 3b. Each tool $t\in\mathcal{T}$ is represented as an action with a clearly defined set of the following three components: (1) Parameters, define the input arguments for the tool, including data types and constraints; (2) Preconditions, denoted as $\text{Pre}(t)$, record logical conditions that must be met before tool execution; (3) Effects, denoted as $\text{Eff}(t)$), evaluate the state changes that occur upon tool execution.
We formalize the operational space of the agent, denoted as $\mathbf{\Omega}$, as a structured hierarchical taxonomy rather than a simple collection of functions, which aligns with the intrinsic multi-level nature of RS tasks. Specifically, $\mathbf{\Omega}$ is organized into two categories: (1) Atomic operators, which handle low-level data manipulation (e.g., geometric correction, projection); (2) Semantic operators, which encapsulate complex analysis tasks (e.g., segmentation, inversion) into high-level function calls. This rigorous taxonomy ensures that the derived tool space $\mathcal{T}\subset\mathbf{\Omega}$ provides the necessary granularity for domain-specific benchmarks. By unifying these heterogeneous functions into a machine-readable library, we establish the fundamental building blocks for constructing our workflow DAGs. We will further detail the concrete instantiation of this toolset, including specific GDAL and deep learning algorithms, in Section 4.1.1.
@@ 3.2.3 Workflow Templates
We designed this hierarchical structure to enable efficient, top-down planning. By allowing the agent to retrieve and instantiate high-level templates for common tasks, it reduces the planning search space and ensures adherence to expert-validated procedures. This structural guidance constrains the generative process of the LLM, directing it to produce plans that are both syntactically correct and compliant with standard RS protocols.
@@ 3.3 Dynamic Workflow Adjustment
One key innovation of the CangLing-KnowFlow framework is its ability to adjust the workflow in response to runtime failures dynamically, thereby overcoming the inflexibility of static plans. This mechanism is the practical realization of the agent’s policy $\mathcal{A}$ from Eq. (1) under conditions of uncertainty. In RS workflows, the source of this uncertainty could be complicated, such as ambiguities in the user’s prompts (e.g., vague spatial definitions), intrinsic data irregularities (e.g., unexpected cloud cover or sensor noise), and unexpected tool behaviors (e.g., non-convergence of neural networks). When a failure occurs for the abovementioned reasons, the agent must compute a new action $a_{t}$ that deviates from the original plan. This is achieved through a structured process composed of three stages: failure detection, a hierarchical repair strategy, and workflow graph manipulations.
@@ 3.3.1 Failure Detection and Diagnosis
The dynamic adjustment process is initiated when the Dynamic Execution Engine detects a failure at a specific vertex $v_{f}\in V$ of the current workflow $W_{current}$. We define a failure as either an explicit error returned by a tool or an output that fails to meet a predefined quality metric. Once detected, the engine immediately halts the execution of dependent downstream nodes and reports the failure state, which becomes part of the history $\mathcal{H}_{t-1}$, to the Orchestrator Agent. The Orchestrator then diagnoses the root cause by analyzing the history in conjunction with the preconditions of the failed tool $t_{f}$, aiming to identify the root cause of the failure.
@@ 3.3.2 Hierarchical Repair Strategy
Once a failure is diagnosed, the Orchestrator employs a hierarchical strategy to generate a repair plan. Specifically, it selects the next action $a_{t}$ by prioritizing the knowledge from past experience $\mathcal{K}_{evo}$ over generalized reasoning.
@@
    Tier 2: LLM-driven Creative Repair. If no relevant solution is found in $\mathcal{K}_{evo}$, the agent pushes the problem to the LLM. It provides the LLM with the full context ($\mathcal{H}_{t-1}$, $W_{current}$) to generate a novel repair plan, which then defines the action $a_{t}$.
@@ 3.3.3 Workflow Graph Manipulation
The repair plan, formalized as the chosen action $a_{t}$, is executed by applying a series of topological transformations to the current workflow DAG $W_{current}$, resulting in an updated workflow structure $W^{\prime}_{current}$. These transformations primarily consist of three distinct graph operations designed to rectify the identified fault. The first transformation, Node Replacement, involves substituting the failed vertex $v_{f}$ with a new vertex $v_{r}$ associated with an alternative tool $t_{r}$, ensuring that all original incoming and outgoing edges of $v_{f}$ are correctly remapped to $v_{r}$. Alternatively, the agent may perform Node Insertion, which introduces a new intermediate vertex $v_{new}$ between a predecessor node $v_{p}$ and the failed node $v_{f}$. This operation structurally modifies the path by severing the direct link $(v_{p},v_{f})$ and establishing a new sequence of edges $(v_{p},v_{new})$ and $(v_{new},v_{f})$. Finally, Parameter Modification allows for in-situ adjustment, where the vertex $v_{f}$ is topologically retained, but the hyperparameters governing its underlying tool $t_{f}$ are recalibrated to resolve the execution error.
@@ 3.4 Experience-driven Memory and Evolution
This module plays a role that endows the CangLing-KnowFlow agent with the ability to learn and evolve from its operational experience, addressing the critical learning gap in existing agentic workflows. It is mainly achieved through a post-task analysis and knowledge generalization process that operates on a repository of detailed execution traces.
@@ 3.4.1 Structure of the Evolutionary Memory
The memory module serves as the persistent storage for the evolutionary knowledge $\mathcal{K}_{evo}$ introduced in Eq. (1). Structurally, it is organized as a database of execution traces. Each trace, $\mathcal{E}$, is a comprehensive log of a completed task, containing the initial user request, the initial workflow DAG $W_{init}$, the full sequence of actions taken, all intermediate data and tool outputs, any failures encountered, the dynamic adjustments performed, and the final outcome. This rich and structured dataset serves as the solid foundation for the learning processes of our agent.
@@ 3.4.2 Knowledge Generalization and Refinement
Success Solidification (Positive Feedback): If a task that involved dynamic adjustments is completed successfully, the final validated workflow $W_{final}$ can then be considered a candidate worth promoting. After the agent analyzes the trace to identify the specific sequence of adjustments that led to success, the successful, dynamically generated workflow is then generalized into a new, reusable template. This template, along with metadata describing the context in which it was effective, will be finally added to the PKB. In this manner, the whole process could continuously enrich the PKB with novel, field-tested procedures.
Failure Attribution (Negative Feedback): In the event of a terminal task failure, the agent performs a failure attribution analysis on the execution trace $\mathcal{E}$. It identifies the specific failure point $v_{f}$, the state of the workflow at that time, and the diagnosed root cause. Based on this analysis, it synthesizes a new Pattern-Action Rule, formally denoted as $r\in\mathcal{K}_{evo}$. This rule represents a conditional logic that maps a specific failure pattern to a recommended repair action:
IF tool(’YOLOvax’) fails on data(type=’SAR’) with error(’low_contrast’)
THEN Apply action(insert_tool(’Speckle_Suppression’))
These generated rules are stored in a dedicated section of the memory module and are prioritized during the “Knowledge-driven Repair” tier of the dynamic adjustment process. This creates a negative feedback loop that allows the agent to learn from its mistakes and provide a robust repair strategy when errors occur, significantly reducing the latency and token costs associated with repetitive LLM querying.
Through the dual mechanism of solidifying successes into templates and distilling failures into repair rules, the CangLing-KnowFlow agent can systematically transform its episodic experiences into durable procedural knowledge ($\mathcal{K}_{evo}$), enabling it to evolve and increase its accuracy over time.
@@ 4 Database and Benchmark
To ground our agent’s capabilities and rigorously evaluate its performance, we constructed two key databases: (1) an internal PKB to provide the agent with domain-specific knowledge and procedural guidance, and (2) a comprehensive evaluation suite, termed KnowFlow-Bench, to assess its workflow orchestration capabilities against expert-defined standards.
@@ 4.1 Procedural Knowledge Base Construction
The PKB is the cornerstone of the CangLing-KnowFlow framework, designed to infuse the agent with expert-level procedural knowledge. This section details the practical construction of the PKB, whose formal structure and mathematical preliminaries were defined in Section 3.2. The construction involved a two-stage process of toolset formalization and high-level workflow curation.
@@ 4.1.1 Toolset Formalization
Deep Learning Tools: We designed a suite of high-level deep learning tools, each tailored to a primary RS analysis task. These tools decouple functional logic from architectural complexity, allowing the system to focus on the transformation of input data into analytical outputs. The suite includes four main categories: semantic_segmentation, object_detection, change_detection, and remote_sensing_inversion. Each tool is defined by a rigorous JSON schema that specifies its functionality, parameters, and input/output requirements, ensuring reliable interpretation and invocation by the agent. For example, the semantic_segmentation tool requires parameters such as image_path, target_classes, and spatial_resolution_m to execute pixel-wise classification.
@@ 4.2.1 The KnowFlow-Bench
Existing benchmarks typically focus on isolated capabilities, for example, ThinkGeo [shabbir2025thinkgeo] assesses step-by-step tool usage, while GeoCode-Bench [hou2025can] centers specifically on geospatial code generation. Consequently, the field lacks a dedicated standard for evaluating an agent’s comprehensive capacity to translate high-level goals into executable workflows.
To bridge this gap, KnowFlow-Bench is built upon the 162 tasks curated from our PKB. For each task, domain experts annotated two distinct, complete, and executable ground-truth workflows, yielding a total of 324 evaluation instances. Unlike the high-level templates in the PKB, each workflow in KnowFlow-Bench consists of a concrete sequence of function calls with fully specified parameters. These “gold standard” workflows represent scientifically valid solution paths given the available toolset, serving as the reference for evaluating both the procedural correctness and final output accuracy of the agent.
@@ 4.2.2 Cross-Benchmark Generalization on ThinkGeo
We also extended our evaluation to the ThinkGeo benchmark [shabbir2025thinkgeo] to test the transferability and generalization of our framework. ThinkGeo is a comprehensive agentic benchmark comprising 436 tasks grounded in real-world RS imagery, encompassing seven major domains such as urban planning and disaster assessment. It provides a suite of 14 predefined tools and evaluates agents based on their ability to follow a ReAct-style reasoning and execution chain. To facilitate this cross-evaluation, we configured the CangLing-KnowFlow agent to align with the ThinkGeo environment. This adaptation entailed extracting the procedural workflow logic from its ReAct-style annotations and restricting the agent’s action space strictly to the 14 tools provided by the benchmark. This set of experiments highlights the robustness of our agent’s planning mechanisms, confirming its effectiveness even when operating with external tool definitions and novel task requirements.